# Understanding Deep Representation Learning via Layerwise Feature Compression and Discrimination

### v3

@@ Appendix B Proofs of Main Results
In this section, we provide formal proofs to show our main result in 1 concerning the behavior of within-class and between-class features. This involves establishing some key properties of the weight matrices of each layer under Assumptions 1 and 2. The rest of this section is organized as follows: In Section B.1, we establish some preliminary results for weight matrices under Assumptions 1 and 2. In Section B.2, we prove that within-class features are compressed across layers at a geometric rate. In Section B.3, we prove that between-class features are discriminated across layers at a linear rate. Finally, we finish the proofs of 1 and 1 in Section B.4 and B.5, respectively.
@@
Before we proceed, let us introduce some further notation that will be used throughout this section. Using (10), $\bm{Y}=\bm{I}_{K}\otimes\bm{1}_{n}^{\top}$, and Assumption 1, we show that the rank of $\bm{W}_{l}$ is at least $K$ for each $l\in[L]$; see Lemma 7. For all $l\in[L-1]$, let
$\bm{W}_{l}=\bm{U}_{l}\bm{\Sigma}_{l}\bm{V}_{l}^{\top}=\begin{bmatrix}\bm{U}_{l,1}&\bm{U}_{l,2}\end{bmatrix}\begin{bmatrix}\bm{\Sigma}_{l,1}&\bm{0}\\ \bm{0}&\bm{\Sigma}_{l,2}\end{bmatrix}\begin{bmatrix}\bm{V}_{l,1}^{\top}\\ \bm{V}_{l,2}^{\top}\end{bmatrix}=\bm{U}_{l,1}\bm{\Sigma}_{l,1}\bm{V}_{l,1}^{\top}+\bm{U}_{l,2}\bm{\Sigma}_{l,2}\bm{V}_{l,2}^{\top},$ (25)
@@
$\bm{W}_{L}=\bm{U}_{L}\bm{\Sigma}_{L}\bm{V}_{L}^{\top}=\bm{U}_{L}\begin{bmatrix}\bm{\Sigma}_{L,1}&\bm{0}\end{bmatrix}\begin{bmatrix}\bm{V}_{L,1}^{\top}\\ \bm{V}_{L,2}^{\top}\end{bmatrix}=\bm{U}_{L}\bm{\Sigma}_{L,1}\bm{V}_{L,1}^{\top},$ (26)
@@
$\bar{\bm{X}}=\left[\bar{\bm{x}}_{1}\ \bar{\bm{x}}_{2}\ \dots\ \bar{\bm{x}}_{K}\right]\otimes\bm{1}_{n}^{\top}\in\mathbb{R}^{d\times N}.$ (28)
Note that we shall use the above notation repeatedly in the following context.
B.1 Properties of Weight Matrices
In this subsection, we establish some properties of weight matrices $\{\bm{W}_{l}\}_{l=1}^{L}$ under Assumptions 1 and 2. Towards this goal, we first prove some inequalities on the spectrum of the data matrix $\bm{X}$ and the distance between $\bm{X}$ and $\bar{\bm{X}}$ under Assumption 1. These inequalities will serve as crucial components in the subsequent analysis of the spectral properties of the weight matrices.
Lemma 5.
@@ B.1 Properties of Weight Matrices
Proof.
Using $\|\bm{A}\|\leq\|\bm{A}\|_{F}$ for all $\bm{A}\in\mathbb{R}^{N\times N}$, we have
$\|\bm{X}^{\top}\bm{X}-\bm{I}_{N}\|^{2}\leq\|\bm{X}^{\top}\bm{X}-\bm{I}_{N}\|_{F}^{2}\leq\sum_{i=1}^{N}\left(\|\bm{x}_{i}\|^{2}-1\right)^{2}+\sum_{i=1}^{N}\sum_{j\neq i}^{N}\left(\bm{x}_{i}^{\top}\bm{x}_{j}\right)^{2}\leq\theta^{2},$ (31)
where the last inequality follows from (9). This, together with Weyl’s inequality, implies $1-\theta\leq\sigma_{\min}(\bm{X}^{\top}\bm{X})\leq\sigma_{\max}(\bm{X}^{\top}\bm{X})\leq 1+\theta$. Therefore, we obtain (29).
Now, we prove (30). For ease of exposition, let $\bm{A}:=(\bm{X}-\bar{\bm{X}})^{\top}(\bm{X}-\bar{\bm{X}})=\bm{X}^{\top}\bm{X}-\bm{X}^{\top}\bar{\bm{X}}-\bar{\bm{X}}^{\top}\bm{X}+\bar{\bm{X}}^{\top}\bar{\bm{X}}$. First, we consider the elements of diagonal blocks of $\bm{A}$. For each $i=(k-1)n+j$ and $j\in[n]$, we compute
@@
Using the similar argument, we compute for each $i=(k-1)n+j$, $i^{\prime}=(k-1)n+j^{\prime}$, and $j\neq j^{\prime}\in[n]$,
@@
Now, we consider the elements of off-diagonal blocks of $\bm{A}$. Similarly, we compute for each $i=(k-1)n+j$, $i^{\prime}=(k^{\prime}-1)n+j^{\prime}$, $k\neq k^{\prime}$, and $j,j^{\prime}\in[n]$,
@@
This, together with (32) and (33), yields that
$\bm{A}=\bm{I}_{N}-\frac{1}{n}I_{K}\otimes\bm{E}_{n}+\bm{\Delta},$ (35)
where $|\delta_{ij}|\leq{4\theta}/{N}$. Therefore, we have
@@
which implies the first inequality in (30).
Moreover, we have
@@
Then, we will bound the above terms in turn. First, noting that $\|\bm{X}\|_{F}^{2}=\sum_{i=1}^{N}\|\bm{x}_{i}\|^{2}$ and using (9) in Assumption 1 yield
@@
This, together with Assumption 1, yields
$K-\theta\leq\|\bar{\bm{X}}\|_{F}^{2}\leq K+\theta.$ (39)
Substituting (37), (38), and (39) into (36) yields the second inequality in (30). ∎
Next, we present a lemma that shows that each singular value of the first $L-1$ weight matrices is exactly the same, and each singular value of the last weight matrix is approximately equal to that of the first $L-1$ weight matrices under the balancedness (11) in Assumption 2.
Lemma 6.
@@
$\bm{\Sigma}_{l+1}=\bm{\Sigma}_{l},\ \forall l\in[L-2],\ \|\bm{\Sigma}_{L}^{\top}\bm{\Sigma}_{L}-\bm{\Sigma}_{l}^{2}\|_{F}\leq\delta,\ \forall l\in[L-1].$ (40)
Proof.
Since $\bm{W}_{l+1}^{\top}\bm{W}_{l+1}=\bm{W}_{l}\bm{W}_{l}^{\top}$ for all $l\in[L-2]$ and (25), we have
$\bm{V}_{l+1}\bm{\Sigma}_{l+1}^{2}\bm{V}_{l+1}^{\top}=\bm{U}_{l}\bm{\Sigma}_{l}^{2}\bm{U}_{l}^{\top}.$ (41)
For a given $l\in[L-2]$, the two sides of the above equation are essentially eigenvalue decomposition of the same matrix. This, together with the fact that $\bm{\Sigma}_{l+1}$ and $\bm{\Sigma}_{l}$ have non-increasing and non-negative diagonal elements by (25), yields $\bm{\Sigma}_{l+1}=\bm{\Sigma}_{l}$. Using $\|\bm{W}_{L}^{\top}\bm{W}_{L}-\bm{W}_{L-1}\bm{W}_{L-1}^{\top}\|_{F}\leq\delta$, (25), and (26), we have
$\|\bm{V}_{L}\bm{\Sigma}_{L}^{\top}\bm{\Sigma}_{L}\bm{V}_{L}^{\top}-\bm{U}_{L-1}\bm{\Sigma}_{L-1}^{2}\bm{U}_{L-1}^{\top}\|_{F}\leq\delta.$ (42)
This, together with [2, Lemma 4], implies $\|\bm{\Sigma}_{L}^{\top}\bm{\Sigma}_{L}-\bm{\Sigma}_{L-1}^{2}\|_{F}\leq\delta$. According to this and $\bm{\Sigma}_{l+1}=\bm{\Sigma}_{l}$ for all $l\in[L-2]$, implies (40) for all $l\in[L-1]$. ∎
Based on the above two lemmas, we are ready to estimate the top $K$ singular values of each weight matrix $\bm{W}_{l}$ for all $l\in[L]$ under Assumptions 1 and 2.
Lemma 7.
@@
Proof.
According to (29) in Lemma 5, we obtain that $\bm{X}$ is of full column rank, and thus $\bm{X}^{+}:=(\bm{X}^{\top}\bm{X})^{-1}\bm{X}^{\top}$ is of full row rank with
$1/\sqrt{1+\theta}\leq\sigma_{\min}(\bm{X}^{+})\leq\sigma_{\max}(\bm{X}^{+})\leq 1/\sqrt{1-\theta}.$ (48)
This, together with $\bm{Y}=\bm{I}_{K}\otimes\bm{1}_{n}^{\top}$, (10), and Lemma 2, yields that
@@
Using this, (11) with (43), and invoking [2, Lemma 6], we obtain $\sigma_{1}(\bm{W}_{l})\leq\left(2\sqrt{n}\right)^{1/L}$ for all $l\in[L]$. Using $\bm{W}_{l+1}^{\top}\bm{W}_{l+1}=\bm{W}_{l}\bm{W}_{l}^{\top}$ for all $l\in[L-2]$ by (11), we compute
$\|\bm{W}_{L:1}^{\top}\bm{W}_{L:1}-\left(\bm{W}_{1}^{\top}\bm{W}_{1}\right)^{L}\|_{F}$ $=\|\bm{W}_{L:1}^{\top}\bm{W}_{L:1}-\bm{W}_{L-1:1}^{\top}\bm{W}_{L-1}\bm{W}_{L-1}^{\top}\bm{W}_{L-1:1}\|_{F}$
$=\|\bm{W}_{L-1:1}^{\top}\left(\bm{W}_{L}^{\top}\bm{W}_{L}-\bm{W}_{L-1}\bm{W}_{L-1}^{\top}\right)\bm{W}_{L-1:1}\|_{F}$
$\leq\|\bm{W}_{L}^{\top}\bm{W}_{L}-\bm{W}_{L-1}\bm{W}_{L-1}^{\top}\|_{F}\prod_{l=1}^{L-1}\|\bm{W}_{l}\|^{2}$
@@
$\sigma_{K}\left(\left(\bm{W}_{1}^{\top}\bm{W}_{1}\right)^{L}\right)$ $\geq\sigma_{K}\left(\bm{W}_{L:1}^{\top}\bm{W}_{L:1}\right)-\|\bm{W}_{L:1}^{\top}\bm{W}_{L:1}-\left(\bm{W}_{1}^{\top}\bm{W}_{1}\right)^{L}\|$
$\geq\frac{n}{1+\theta}-\|\bm{W}_{L:1}^{\top}\bm{W}_{L:1}-\left(\bm{W}_{1}^{\top}\bm{W}_{1}\right)^{L}\|_{F}\geq\left(\frac{1}{1+\theta}-\frac{2}{15L^{2}}\right)n,$
which implies $\sigma_{K}(\bm{W}_{1})\geq\left(\left(\frac{1}{1+\theta}-\frac{2}{15L^{2}}\right)n\right)^{1/2L}$. This, together with Lemma 6, further implies
@@
$\sigma_{K}\left(\bm{W}_{L}^{\top}\bm{W}_{L}\right)$ $\geq\sigma_{K}\left(\bm{W}_{L-1}^{\top}\bm{W}_{L-1}\right)-\|\bm{W}_{L}^{\top}\bm{W}_{L}-\bm{W}_{L-1}\bm{W}_{L-1}^{\top}\|$
@@
Next, we are devoted to proving (46) and (47). It follows from (B.1) that
$\|\bm{W}_{L:1}^{\top}\bm{W}_{L:1}-\left(\bm{W}_{1}^{\top}\bm{W}_{1}\right)^{L}\|_{F}\leq 4n\delta.$ (53)
@@
$\sigma_{1}\left(\left(\bm{W}_{1}^{\top}\bm{W}_{1}\right)^{L}\right)$ $\leq\sigma_{1}\left(\bm{W}_{L:1}^{\top}\bm{W}_{L:1}\right)+\|\bm{W}_{L:1}^{\top}\bm{W}_{L:1}-\left(\bm{W}_{1}^{\top}\bm{W}_{1}\right)^{L}\|\leq\left(\frac{1}{1-\theta}+4\delta\right)n,$
$\sigma_{K}\left(\left(\bm{W}_{1}^{\top}\bm{W}_{1}\right)^{L}\right)$ $\geq\sigma_{K}\left(\bm{W}_{L:1}^{\top}\bm{W}_{L:1}\right)-\|\bm{W}_{L:1}^{\top}\bm{W}_{L:1}-\left(\bm{W}_{1}^{\top}\bm{W}_{1}\right)^{L}\|\geq\left(\frac{1}{1+\theta}-4\delta\right)n.$
@@
$\sigma_{1}\left(\bm{W}_{L}^{\top}\bm{W}_{L}\right)$ $\leq\sigma_{1}\left(\bm{W}_{L-1}^{\top}\bm{W}_{L-1}\right)+\|\bm{W}_{L}^{\top}\bm{W}_{L}-\bm{W}_{L-1}\bm{W}_{L-1}^{\top}\|\leq\left(\frac{1}{1-\theta}+4\delta\right)^{1/L}n^{1/L}+\delta,$
$\sigma_{K}\left(\bm{W}_{L}^{\top}\bm{W}_{L}\right)$ $\geq\sigma_{K}\left(\bm{W}_{L-1}^{\top}\bm{W}_{L-1}\right)-\|\bm{W}_{L}^{\top}\bm{W}_{L}-\bm{W}_{L-1}\bm{W}_{L-1}^{\top}\|\geq\left(\frac{1}{1-\theta}+4\delta\right)^{1/L}n^{1/L}-\delta,$
which directly implies (47). ∎
Using this lemma and $\varepsilon\in(0,1)$, we conclude that the leading $K$ singular values of $\bm{W}_{l}$ for all $l\in[L-1]$ are well separated from the remaining $d-K$ singular values according to (12). Consequently, we can handle them with the associated singular vectors separately. Since $\bm{\Sigma}_{l+1}=\bm{\Sigma}_{l}$ for all $l\in[L-2]$ according to Lemma 6, suppose that $\bm{\Sigma}_{l,1}$ and $\bm{\Sigma}_{l,2}$ have distinct $p\geq 1$ and $q\geq 1$ singular values for all $l\in[L-1]$, respectively. Let $\sigma_{1}>\dots>\sigma_{p}>\sigma_{p+1}>\dots>\sigma_{p+q}$ be the distinct singular values with the corresponding multiplicities $h_{1},h_{2},\dots,h_{p+q}\in\mathbb{N}$. In particular, we have $\sum_{i=1}^{p}h_{i}=K$, $\sum_{i=p+1}^{p+q}h_{i}=d-K$, and $\sigma_{p+1}\leq\varepsilon$ due to (12). Then, we can write
@@
Lemma 8.
Suppose that Assumptions 1 and 2 hold. Then, the following statements hold:
@@
[…]
Proof.
(i) Using the first equation in (40), (41), (54), and (55), we obtain
$\sum_{i=1}^{p+q}\sigma_{i}^{2}\bm{V}_{l+1}^{(i)}\bm{V}_{l+1}^{(i)^{\top}}=\sum_{i=1}^{p+q}\sigma_{i}^{2}\bm{U}_{l}^{(i)}\bm{U}_{l}^{(i)^{\top}}$

### v4

@@ Appendix A Proofs of Main Results in Section 3
In this section, we provide formal proofs to show our main result in Theorem 4 concerning the behavior of within-class and between-class features. This involves establishing some key properties of the weight matrices of each layer under Assumptions 2 and 3. The rest of this section is organized as follows: In Section A.1, we establish some preliminary results for weight matrices under Assumptions 2 and 3. In Section A.2, we prove that within-class features are compressed across layers at a geometric rate. In Section A.3, we prove that between-class features are discriminated across layers at a linear rate. Finally, we finish the proofs of Theorem 4 and Proposition 5 in Section A.4 and A.5, respectively.
@@
Before we proceed, let us introduce some further notation that will be used throughout this section. Using (10), $\bm{Y}=\bm{I}_{K}\otimes\bm{1}_{n}^{T}$, and Assumption 2, we show that the rank of $\bm{W}_{l}$ is at least $K$ for each $l\in[L]$; see Lemma 8. For all $l\in[L-1]$, let
$\bm{W}_{l}=\bm{U}_{l}\bm{\Sigma}_{l}\bm{V}_{l}^{T}=\begin{bmatrix}\bm{U}_{l,1}&\bm{U}_{l,2}\end{bmatrix}\begin{bmatrix}\bm{\Sigma}_{l,1}&\bm{0}\\ \bm{0}&\bm{\Sigma}_{l,2}\end{bmatrix}\begin{bmatrix}\bm{V}_{l,1}^{T}\\ \bm{V}_{l,2}^{T}\end{bmatrix}=\bm{U}_{l,1}\bm{\Sigma}_{l,1}\bm{V}_{l,1}^{T}+\bm{U}_{l,2}\bm{\Sigma}_{l,2}\bm{V}_{l,2}^{T},$ (21)
@@
$\bm{W}_{L}=\bm{U}_{L}\bm{\Sigma}_{L}\bm{V}_{L}^{T}=\bm{U}_{L}\begin{bmatrix}\bm{\Sigma}_{L,1}&\bm{0}\end{bmatrix}\begin{bmatrix}\bm{V}_{L,1}^{T}\\ \bm{V}_{L,2}^{T}\end{bmatrix}=\bm{U}_{L}\bm{\Sigma}_{L,1}\bm{V}_{L,1}^{T},$ (22)
@@
$\bar{\bm{X}}=\left[\bar{\bm{x}}_{1}\ \bar{\bm{x}}_{2}\ \dots\ \bar{\bm{x}}_{K}\right]\otimes\bm{1}_{n}^{T}\in\mathbb{R}^{d\times N}.$ (24)
Note that the above notation will be used consistently throughout the rest of this paper.
A.1 Properties of Weight Matrices
In this subsection, we establish some properties of weight matrices $\{\bm{W}_{l}\}_{l=1}^{L}$ under Assumptions 2 and 3. Towards this goal, we first prove some inequalities on the spectrum of the data matrix $\bm{X}$ and the distance between $\bm{X}$ and $\bar{\bm{X}}$ under Assumption 2. These inequalities will serve as crucial components in the subsequent analysis of the spectral properties of the weight matrices.
Lemma 6
@@ A.1 Properties of Weight Matrices
Proof Using $\|\bm{A}\|\leq\|\bm{A}\|_{F}$ for all $\bm{A}\in\mathbb{R}^{N\times N}$, we have
$\|\bm{X}^{T}\bm{X}-\bm{I}_{N}\|^{2}\leq\|\bm{X}^{T}\bm{X}-\bm{I}_{N}\|_{F}^{2}\leq\sum_{i=1}^{N}\left(\|\bm{x}_{i}\|^{2}-1\right)^{2}+\sum_{i=1}^{N}\sum_{j\neq i}^{N}\left(\bm{x}_{i}^{T}\bm{x}_{j}\right)^{2}\leq\theta^{2},$ (27)
where the last inequality follows from (9). This, together with Weyl’s inequality, implies $1-\theta\leq\sigma_{\min}(\bm{X}^{T}\bm{X})\leq\sigma_{\max}(\bm{X}^{T}\bm{X})\leq 1+\theta$. Therefore, we obtain (25).
Now, we prove (26). For ease of exposition, let $\bm{A}:=(\bm{X}-\bar{\bm{X}})^{T}(\bm{X}-\bar{\bm{X}})=\bm{X}^{T}\bm{X}-\bm{X}^{T}\bar{\bm{X}}-\bar{\bm{X}}^{T}\bm{X}+\bar{\bm{X}}^{T}\bar{\bm{X}}$. First, we consider the elements of diagonal blocks of $\bm{A}$. For each i=(k1)n+ji=(k-1)n+j and $j\in[n]$, we compute
@@
Using the similar argument, we compute for each i=(k1)n+ji=(k-1)n+j, $i^{\prime}=(k-1)n+j^{\prime}$, and $j\neq j^{\prime}\in[n]$,
@@
Now, we consider the elements of off-diagonal blocks of $\bm{A}$. Similarly, we compute for each i=(k1)n+ji=(k-1)n+j, $i^{\prime}=(k^{\prime}-1)n+j^{\prime}$, $k\neq k^{\prime}$, and $j,j^{\prime}\in[n]$,
@@
This, together with (28) and (29), yields $\bm{A}=\bm{I}_{N}-\frac{1}{n}\bm{I}_{K}\otimes\bm{E}_{n}+\bm{\Delta}$, where $|\delta_{ij}|\leq{4\theta}/{N}$. Therefore, we have
@@
which implies the first inequality in (26). Moreover, we have
@@
Then, we bound the above terms in turn. First, noting that $\|\bm{X}\|_{F}^{2}=\sum_{i=1}^{N}\|\bm{x}_{i}\|^{2}$ and using (9) in Assumption 2 yield
@@
This, together with Assumption 2, yields $K-\theta\leq\|\bar{\bm{X}}\|_{F}^{2}\leq K+\theta$. Substituting this, (32), and (33) into (31) yields the second inequality in (26).
Next, we present a lemma that shows that each singular value of the first L1L-1 weight matrices is exactly the same, and each singular value of the last weight matrix is approximately equal to that of the first L1L-1 weight matrices under the balancedness (11) in Assumption 3.
Lemma 7
@@
$\bm{\Sigma}_{l+1}=\bm{\Sigma}_{l},\ \forall l\in[L-2],\ \|\bm{\Sigma}_{L}^{T}\bm{\Sigma}_{L}-\bm{\Sigma}_{l}^{2}\|_{F}\leq\delta,\ \forall l\in[L-1].$ (34)
Proof Since $\bm{W}_{l+1}^{T}\bm{W}_{l+1}=\bm{W}_{l}\bm{W}_{l}^{T}$ for all $l\in[L-2]$ and (21), we have
$\bm{V}_{l+1}\bm{\Sigma}_{l+1}^{2}\bm{V}_{l+1}^{T}=\bm{U}_{l}\bm{\Sigma}_{l}^{2}\bm{U}_{l}^{T}.$ (35)
For a given $l\in[L-2]$, the two sides of the above equation are essentially eigenvalue decomposition of the same matrix. This, together with the fact that $\bm{\Sigma}_{l+1}$ and $\bm{\Sigma}_{l}$ have non-increasing and non-negative diagonal elements by (21), yields $\bm{\Sigma}_{l+1}=\bm{\Sigma}_{l}$. Using $\|\bm{W}_{L}^{T}\bm{W}_{L}-\bm{W}_{L-1}\bm{W}_{L-1}^{T}\|_{F}\leq\delta$, (21), and (22), we have
$\|\bm{V}_{L}\bm{\Sigma}_{L}^{T}\bm{\Sigma}_{L}\bm{V}_{L}^{T}-\bm{U}_{L-1}\bm{\Sigma}_{L-1}^{2}\bm{U}_{L-1}^{T}\|_{F}\leq\delta.$ (36)
This, together with (arora2018convergence, Lemma 4), implies $\|\bm{\Sigma}_{L}^{T}\bm{\Sigma}_{L}-\bm{\Sigma}_{L-1}^{2}\|_{F}\leq\delta$. According to this and $\bm{\Sigma}_{l+1}=\bm{\Sigma}_{l}$ for all $l\in[L-2]$, implies (34) for all $l\in[L-1]$.
Based on the above two lemmas, we are ready to estimate the top $K$ singular values of each weight matrix $\bm{W}_{l}$ for all $l\in[L]$ under Assumptions 2 and 3.
Lemma 8
@@
Proof According to (25) in Lemma 6, we obtain that $\bm{X}$ is of full column rank, and thus $\bm{X}^{+}:=(\bm{X}^{T}\bm{X})^{-1}\bm{X}^{T}$ is of full row rank with $1/\sqrt{1+\theta}\leq\sigma_{\min}(\bm{X}^{+})\leq\sigma_{\max}(\bm{X}^{+})\leq 1/\sqrt{1-\theta}$. This, together with $\bm{Y}=\bm{I}_{K}\otimes\bm{1}_{n}^{T}$, (10), and Lemma 15, yields
@@
Using this, (11) with (37), and invoking (arora2018convergence, Lemma 6), we obtain $\sigma_{1}(\bm{W}_{l})\leq\left(2\sqrt{n}\right)^{1/L}$ for all $l\in[L]$. Using $\bm{W}_{l+1}^{T}\bm{W}_{l+1}=\bm{W}_{l}\bm{W}_{l}^{T}$ for all $l\in[L-2]$ by (11), we compute
$\|\bm{W}_{L:1}^{T}\bm{W}_{L:1}-\left(\bm{W}_{1}^{T}\bm{W}_{1}\right)^{L}\|_{F}$ $=\|\bm{W}_{L:1}^{T}\bm{W}_{L:1}-\bm{W}_{L-1:1}^{T}\bm{W}_{L-1}\bm{W}_{L-1}^{T}\bm{W}_{L-1:1}\|_{F}$
$=\|\bm{W}_{L-1:1}^{T}\left(\bm{W}_{L}^{T}\bm{W}_{L}-\bm{W}_{L-1}\bm{W}_{L-1}^{T}\right)\bm{W}_{L-1:1}\|_{F}$
$\leq\|\bm{W}_{L}^{T}\bm{W}_{L}-\bm{W}_{L-1}\bm{W}_{L-1}^{T}\|_{F}\prod_{l=1}^{L-1}\|\bm{W}_{l}\|^{2}$
@@
$\sigma_{K}\left(\left(\bm{W}_{1}^{T}\bm{W}_{1}\right)^{L}\right)$ $\geq\sigma_{K}\left(\bm{W}_{L:1}^{T}\bm{W}_{L:1}\right)-\|\bm{W}_{L:1}^{T}\bm{W}_{L:1}-\left(\bm{W}_{1}^{T}\bm{W}_{1}\right)^{L}\|$
$\geq\frac{n}{1+\theta}-\|\bm{W}_{L:1}^{T}\bm{W}_{L:1}-\left(\bm{W}_{1}^{T}\bm{W}_{1}\right)^{L}\|_{F}\geq\left(\frac{1}{1+\theta}-\frac{2}{15L^{2}}\right)n,$
which implies $\sigma_{K}(\bm{W}_{1})\geq\left(\left(\frac{1}{1+\theta}-\frac{2}{15L^{2}}\right)n\right)^{1/2L}$. This, together with Lemma 7, implies
@@
$\sigma_{K}\left(\bm{W}_{L}^{T}\bm{W}_{L}\right)$ $\geq\sigma_{K}\left(\bm{W}_{L-1}^{T}\bm{W}_{L-1}\right)-\|\bm{W}_{L}^{T}\bm{W}_{L}-\bm{W}_{L-1}\bm{W}_{L-1}^{T}\|$
@@
Next, we are devoted to proving (40) and (41). It follows from (A.1) that
$\|\bm{W}_{L:1}^{T}\bm{W}_{L:1}-\left(\bm{W}_{1}^{T}\bm{W}_{1}\right)^{L}\|_{F}\leq 4n\delta.$ (46)
@@
$\sigma_{1}\left(\left(\bm{W}_{1}^{T}\bm{W}_{1}\right)^{L}\right)$ $\leq\sigma_{1}\left(\bm{W}_{L:1}^{T}\bm{W}_{L:1}\right)+\|\bm{W}_{L:1}^{T}\bm{W}_{L:1}-\left(\bm{W}_{1}^{T}\bm{W}_{1}\right)^{L}\|\leq\left(\frac{1}{1-\theta}+4\delta\right)n,$
$\sigma_{K}\left(\left(\bm{W}_{1}^{T}\bm{W}_{1}\right)^{L}\right)$ $\geq\sigma_{K}\left(\bm{W}_{L:1}^{T}\bm{W}_{L:1}\right)-\|\bm{W}_{L:1}^{T}\bm{W}_{L:1}-\left(\bm{W}_{1}^{T}\bm{W}_{1}\right)^{L}\|\geq\left(\frac{1}{1+\theta}-4\delta\right)n.$
@@
$\sigma_{1}\left(\bm{W}_{L}^{T}\bm{W}_{L}\right)$ $\leq\sigma_{1}\left(\bm{W}_{L-1}^{T}\bm{W}_{L-1}\right)+\|\bm{W}_{L}^{T}\bm{W}_{L}-\bm{W}_{L-1}\bm{W}_{L-1}^{T}\|\leq\left(\frac{1}{1-\theta}+4\delta\right)^{1/L}n^{1/L}+\delta,$
$\sigma_{K}\left(\bm{W}_{L}^{T}\bm{W}_{L}\right)$ $\geq\sigma_{K}\left(\bm{W}_{L-1}^{T}\bm{W}_{L-1}\right)-\|\bm{W}_{L}^{T}\bm{W}_{L}-\bm{W}_{L-1}\bm{W}_{L-1}^{T}\|\geq\left(\frac{1}{1-\theta}+4\delta\right)^{1/L}n^{1/L}-\delta,$
which directly implies (41).
Using this lemma and $\varepsilon\in(0,1)$, we conclude that the leading $K$ singular values of $\bm{W}_{l}$ for all $l\in[L-1]$ are well separated from the remaining dKd-K singular values according to (12). Consequently, we can handle them with the associated singular vectors separately. Since $\bm{\Sigma}_{l+1}=\bm{\Sigma}_{l}$ for all $l\in[L-2]$ according to Lemma 7, suppose that $\bm{\Sigma}_{l,1}$ and $\bm{\Sigma}_{l,2}$ have distinct $p\geq 1$ and $q\geq 1$ singular values for all $l\in[L-1]$, respectively. Let $\sigma_{1}>\dots>\sigma_{p}>\sigma_{p+1}>\dots>\sigma_{p+q}$ be the distinct singular values with the corresponding multiplicities $h_{1},h_{2},\dots,h_{p+q}\in\mathbb{N}$. In particular, we have $\sum_{i=1}^{p}h_{i}=K$, $\sum_{i=p+1}^{p+q}h_{i}=d-K$, and $\sigma_{p+1}\leq\varepsilon$ due to (12). Then, we can write
@@
Lemma 9
Suppose that Assumptions 2 and 3 hold. Then, the following statements hold:
@@
$\|\bm{V}_{L,1}^{T}\bm{U}_{L-1,2}\|_{F}\leq\frac{2\sqrt{\delta}\sqrt[4]{K}}{n^{1/2L}},\ \sigma_{\min}\left(\bm{V}_{L,1}^{T}\bm{U}_{L-1,1}\right)\geq 1-\frac{2\sqrt{\delta}\sqrt[4]{K}}{n^{1/2L}}.$ (50)
Proof (i) Using the first equation in (34), (35), (47), and (48), we obtain
$\sum_{i=1}^{p+q}\sigma_{i}^{2}\bm{V}_{l+1}^{(i)}\bm{V}_{l+1}^{(i)^{T}}=\sum_{i=1}^{p+q}\sigma_{i}^{2}\bm{U}_{l}^{(i)}\bm{U}_{l}^{(i)^{T}}$